Hi, can you help me with solving Part B? I provided part A because I think you need the answer for it to get part B.

In a microwave oven, electrons describe circular motion in a magnetic field within a special tube called a magnetron; as you'll learn in Chapter 29, the electrons' motion results in the production of micowaves. The electrons circle at a frequency of 2.25 GHz . The magnetron can accommodate electron orbits with maximum diameter  2.72 mm .

Part A

What's the magnetic field strength? Got the answer as 80.5 mT

Part B

What's the electrons' energy in eV.

K=? eV
Answer #1

Ans:-f = 2.25 GHZ ; w =2πf = 2π*2.25GHz d= 2.75*10^-3m ; r =1.375*10^-3m

b)
K = (1/2)mv²
K = (1/2)mω²r²
K = (1/2)(9.11*10^-31 kg)[(2.25*10^9 s^-1)(2π)]²(1.375*10^-3 m )²
K = 1.72*10^-16J

K=1074.63ev